# SAMix: Calibrated and Accurate Continual Learning via Sphere-Adaptive Mixup and Neural Collapse

### v1

@@ 1 Introduction
To achieve human-like lifelong learning, deep neural networks (DNNs) must support continual learning (CL), the ability to acquire new knowledge incrementally while retaining previously learned information. However, a major challenge is catastrophic forgetting (McCloskey and Cohen 1989; Robins 1995), where new learning disrupts old knowledge. In response, CL has rapidly evolved with numerous approaches (Rusu et al. 2016; Cha, Lee, and Shin 2021; Pourcel, Vu, and French 2022; Wen et al. 2024; Tang et al. 2024; Dang et al. 2025), primarily aimed at balancing adaptation to new data (plasticity) and retention of past knowledge (stability). Compared to joint training (Yoon et al. 2018; Buzzega et al. 2020; Tiwari et al. 2022), decoupling representation and classifier learning has shown superior performance in both continual supervised (Cha, Lee, and Shin 2021; Wen et al. 2024; Li et al. 2024a; Dang et al. 2025) and self-supervised learning (Fini et al. 2022; Cha, Cho, and Moon 2024; Madaan et al. 2022; Mushtaq et al. 2024). Among these, mixup (Zhang et al. 2018) has emerged as a simple yet effective technique, primarily applied in continual self-supervised learning (Madaan et al. 2022; Mushtaq et al. 2024) to promote smoother task transitions and increase sample diversity, especially for old samples in buffers. Beyond reducing forgetting, mixup also acts as a calibration method, aligning predicted confidence scores with actual accuracy. In DNNs, calibration reflects prediction reliability - a crucial factor for safety-sensitive applications that often require CL, such as medical diagnosis (Jiang et al. 2012) and autonomous driving (Grigorescu et al. 2020), where unreliable predictions can pose serious risks. However, the role of mixup in mitigating overconfidence and enhancing CL performance remains underexplored.
@@
Neural collapse (NC), where last-layer features converge to their class means during training, has attracted significant attention in DNNs across both static and dynamic (continual) learning scenarios (Lu and Steinerberger 2020; Fang et al. 2021; Ji et al. 2022; Zhou et al. 2022; Fisher, Meng, and Papyan 2024). Most NC-based methods rely on the dot-regression loss (DR) (Yang et al. 2022), which enforces feature alignment with fixed class prototypes to enhance class separation and reduce misalignment between features and classifiers. Recently, focal neural collapse contrastive (FNC2) (Dang et al. 2025) loss was introduced to leverage NC more flexibly in CL. Unlike DR, which rigidly aligns features with class prototypes, FNC2 optimizes feature relationships, enhancing intra-class diversity and improving generalization. While NC-based methods effectively improve accuracy and mitigate forgetting, their impact on calibration remains unexplored, as does the effect of calibration in NC-based methods on overall model performance. Indeed, as shown in (Li et al. 2024b; Hwang, Kim, and Whang 2025), beyond performance, CL models must also produce well-calibrated predictions. However, existing attempts at calibration in CL remain relatively simple and leave much room for improvement.
@@
    We propose Sphere-Adaptive Mixup (SAMix), a simple yet effective method to improve calibration, mitigate forgetting, and boost accuracy in NC-based CL. SAMix leverages spherical linear interpolation of fixed ETF (Equiangular Tight Frame) prototypes to generate adaptive mixed prototypes, enhancing representation learning and smoothing decision boundaries.  [^1] We note that (Li et al. 2024b; Hwang, Kim, and Whang 2025) aims to calibrate CL models for more reliable predictions, whereas our work goes further by proposing a method that enhances calibration and also mitigates forgetting, and improves accuracy. We also introduce new metrics to evaluate calibration and overconfidence in CL, using them to assess the reliability of NC-based methods.
@@
    We provide theoretical analyses to compare different NC-based methods for handling mixed samples and validate our findings through empirical results.
@@ 2.1 Continual Learning
CL algorithms can be categorized into two main paradigms: pretrained-based and training-from-scratch. Pretrained methods (Wang et al. 2022b, a; Smith et al. 2023; Liu and Yang 2025) leverage large datasets for general representations, while training-from-scratch approaches suit resource-limited settings. These latter algorithms can be grouped into: Architecture-based methods (Rusu et al. 2016; Yoon et al. 2018; Li et al. 2019), which allocate task-specific parameters or expand networks; Regularization-based methods (Zenke, Poole, and Ganguli 2017; Jung et al. 2020; Fini et al. 2022; Park et al. 2019; Tang et al. 2024; Dang et al. 2025), which constrain weight updates to preserve past knowledge; and Rehearsal-based or replay-based methods (Madaan et al. 2022; Buzzega et al. 2020; Cha, Lee, and Shin 2021; Wen et al. 2024; Mushtaq et al. 2024), which replay stored data or use generative models.
A primary challenge in CL is the plasticity-stability dilemma - balancing the integration of new knowledge (plasticity) with the retention of past information (stability). Regularization-based methods often address this via knowledge distillation (Hinton, Vinyals, and Dean 2015), which preserves prior knowledge during training. This may involve intermediate features (Hou et al. 2019) or final outputs (Fini et al. 2022; Cha, Cho, and Moon 2024). Recent advances extend distillation to relational knowledge, such as instance-wise relation distillation (IRD) in Co2L (Cha, Lee, and Shin 2021) and prototype-instance relation distillation (PRD) in CCLIS (Li et al. 2024a). Building on these, (Dang et al. 2025) proposes hardness-softness distillation (HSD), combining IRD and a variant of PRD for fixed prototypes, achieving superior performance. In this work, we adopt HSD to preserve past knowledge through relational consistency with fixed prototypes.
@@ 2.2 Mixup
Mixup (Zhang et al. 2018) is a data augmentation strategy based on the principle of vicinal risk minimization (Chapelle et al. 2000), which interpolates linearly between example pairs and their corresponding labels. The equation of mixup is defined as:
@@
where $(\mathbf{x}_{i},\mathbf{y}_{i})$, $(\mathbf{x}_{j},\mathbf{y}_{j})$ are two random pairs of samples and one-hot labels drawn from the training data, and $\lambda\in[0,1]\sim\text{Beta}(\alpha,\alpha)$, $\alpha\in(0,\infty)$. With simplicity and ease of deployment, mixup has emerged as a well-known method, used widely in both supervised learning (Guo, Mao, and Zhang 2019; Verma et al. 2019) and unsupervised learning (Kim et al. 2020; Ren et al. 2022; Shen et al. 2022) to prevent overfitting and to improve generalization. Moreover, (Thulasidasan et al. 2019; Noh et al. 2023) show that mixup-trained DNNs are significantly better calibrated, with softmax scores closely reflecting actual correctness likelihood.
Mixup in CL. Mixup has recently been used in CL by combining current task samples with random memory instances to promote linearity across tasks. Most studies focus on self-supervised settings (Madaan et al. 2022; Mushtaq et al. 2024), where mixup serves both as an augmentation technique to increase the number of training samples, especially old ones in memory, and as a regularizer to distill knowledge and reduce forgetting.
Remarks. Like traditional mixup and its CL adaptations, we generates new samples via linear interpolation. However, instead of mixing labels, we interpolate class prototypes on a hypersphere. While mixup benefits in CL, its impact on calibration remains underexplored. This work bridges this gap by showing how our method improves calibration, reduces overconfidence, and still boosts accuracy.
@@ 2.3 Neural Collapse
Neural collapse (NC) (Papyan, Han, and Donoho 2020) is a phenomenon observed during the final stage of DNN training, where last-layer features and classifier form a structured geometric pattern on a balanced dataset, aligning with a simplex Equiangular Tight Frame (ETF). Building on this foundation, subsequent studies have proven NC is globally optimal in balanced training with cross-entropy (Lu and Steinerberger 2020; Zhu et al. 2021; Ji et al. 2022) and mean squared error (Poggio and Liao 2020; Han, Papyan, and Donoho 2022; Zhou et al. 2022) losses. In addition, (Yang et al. 2022) try to induce NC by fixing the classifier under imbalanced training, addressing the issue of misalignment between features and classifier.
NC and Mixup. A recent study (Fisher, Meng, and Papyan 2024) shows that mixup induces a distinct geometric structure. In cross-entropy classifiers with one-hot labels, same-class mixed samples align with a simplex ETF, while cross-class ones form smooth channels along decision boundaries, leading to smoother separation. This effect persists even when classifier weights are fixed as a simplex ETF, enhancing both generalization and calibration.
Inducing NC for CL. Inspired by (Galanti, Gyorgy, and Hutter 2022), which showed that NC persists under transfer to new samples or classes, many CL studies (Yang et al. 2023a, b) leverage NC via DR loss (Yang et al. 2022) to reduce forgetting. These methods fix a simplex ETF as classifier prototypes across tasks and align them with sample features. Meanwhile, contrastive learning has gained traction in CL, with remarkable performance from models like Co2L (Cha, Lee, and Shin 2021), CILA (Wen et al. 2024), and CCLIS (Li et al. 2024a). (Dang et al. 2025) introduces FNC2, a contrastive loss that explores NC in CL. Here, we employ both FNC2 and DR losses in our analysis.
@@ 3.2 Overall Architecture
Our method is illustrated in Figure 1. We adopt a two-stage learning framework that enhances invariant representations across tasks. As in prior work (Cha, Lee, and Shin 2021; Wen et al. 2024; Dang et al. 2025), we incorporate two main losses: (1) plasticity loss $\mathcal{L}_{plas}$ to learn new knowledge and (2) stability loss $\mathcal{L}_{stab}$ to distill past information. We predefine a set of fixed, non-learnable class prototypes $\mathbf{P}=\{\mathbf{p_{i}}\}_{i=1}^{K}$, structured as a simplex ETF. These prototypes are shared across tasks and are used in both loss terms during training. To encourage smoother and more linear transitions between tasks, we introduce SAMix, an adaptive mixup strategy tailored for NC-based approaches. In replay settings, a memory buffer $\mathcal{M}$ is maintained using reservoir sampling (Vitter 1985) to store past samples.
Specifically, during training, each batch consists of samples from the current dataset $\mathcal{D}_{t}$ and the buffer $\mathcal{M}$. These are augmented and randomly selected and mixed along with their corresponding prototypes. All views (original and mixed) are then fed into the current model $f^{t}$ for computing $\mathcal{L}_{plas}$, and into the frozen past model $f^{t-1}$ for $\mathcal{L}_{stab}$. Mixed prototypes are used solely for $\mathcal{L}_{plas}$ while predefined fixed prototypes are involved in both losses. An overview of the full algorithm is provided in the Appendix.

### v2

@@ Abstract
keywords:
Continual Learning, Neural Collapse, Model Calibration, Mixup
@@ 1 Introduction
To achieve human-like lifelong learning, deep neural networks (DNNs) must support continual learning (CL), the ability to acquire new knowledge incrementally while retaining previously learned information. However, a major challenge is catastrophic forgetting [mccloskey_1989, robins1995_catastrophic_fr], where new learning disrupts old knowledge. In response, CL has rapidly evolved with numerous approaches [rusu_2016, co2l, dsdm_eccv_2022, wen2024provablecontrastivecontinuallearning, tang_kaizen_wacv_2024, anh_wacv_2025], primarily aimed at balancing adaptation to new data (plasticity) and retention of past knowledge (stability). Compared to joint training [yoon_2018, buzzega2020darkexperiencegeneralcontinual, tiwari2022gcrgradientcoresetbased], decoupling representation and classifier learning has shown superior performance in both continual supervised [co2l, wen2024provablecontrastivecontinuallearning, li_cl_important_sampling_2024, anh_wacv_2025] and self-supervised learning [fini2022cassle, pseudo_negative_icml_2024, madaan_2022_lump, cromomix_mushtaq_eccv_2024]. Among these, mixup [mixup_2018] has emerged as a simple yet effective technique, primarily applied in continual self-supervised learning [madaan_2022_lump, cromomix_mushtaq_eccv_2024] to promote smoother task transitions and increase sample diversity, especially for old samples in buffers. Beyond reducing forgetting, mixup also acts as a calibration method, aligning predicted confidence scores with actual accuracy, as demonstrated in standard supervised learning and further extended to static imbalanced and long-tailed settings by ReMix [remix_2021]. In DNNs, calibration reflects prediction reliability - a crucial factor for safety-sensitive applications that often require CL, such as medical diagnosis [medical_xiaoqian_2012] and autonomous driving [autonomous_driving_2020], where unreliable predictions can pose serious risks. Moreover, recent analyses in class-imbalanced settings indicate that the reliability of predicted probabilities is highly sensitive to distributional assumptions and resampling-based class rebalancing mechanisms [PICCININNI2024104666, carvalhoResamplingApproachesHandle2025]. Since CL naturally induces class imbalance between current and past tasks, this sensitivity poses additional challenges for maintaining reliable calibration under sequential data and evolving representations. However, the role of mixup in mitigating overconfidence and enhancing CL performance remains underexplored.
@@
Neural collapse (NC), where last-layer features converge to their class means during training, has attracted significant attention in DNNs across both static and dynamic (continual) learning scenarios [lu2021neuralcollapsecrossentropyloss, fang_pnas_2021, wenlong_ji_iclr_2022, zhou_icml_2022, fisher2024pushingboundariesmixupsinfluence]. Most NC-based methods rely on the dot-regression loss (DR) [yang_induce_nc_2022], which enforces feature alignment with fixed class prototypes to enhance class separation and reduce misalignment between features and classifiers. Recently, focal neural collapse contrastive (FNC2) [anh_wacv_2025] loss was introduced to leverage NC more flexibly in CL. Unlike DR, which rigidly aligns features with class prototypes, FNC2 optimizes feature relationships, enhancing intra-class diversity and improving generalization. While NC-based methods effectively improve accuracy and mitigate forgetting, their impact on calibration remains unexplored, as does the effect of calibration in NC-based methods on overall model performance. Indeed, recent studies have highlighted the importance of calibration in CL beyond accuracy alone [li2024calibrationcontinuallearningmodels, hwang2025tciltemperaturescalingusing]. In particular, [li2024calibrationcontinuallearningmodels] shows that CL models are not naturally calibrated under non-stationary training compared to joint training, highlighting that forgetting mitigation alone does not ensure calibration. However, existing calibration approaches in CL remain limited in scope, primarily relying on post-hoc, global logit-scaling strategies and typically assuming uniform calibration across tasks. As a result, they do not explicitly account for key CL-specific challenges such as task-wise distribution shift, class imbalance between old and new tasks, or representation drift as learning progresses, leaving substantial room for improvement.
@@
    We propose Sphere-Adaptive Mixup (SAMix), a simple yet effective method to improve calibration, mitigate forgetting, and boost accuracy in NC-based CL. SAMix leverages spherical linear interpolation (Slerp) of fixed ETF (Equiangular Tight Frame) prototypes to generate adaptive mixed prototypes, enhancing representation learning and smoothing decision boundaries.  [^1] We note that [li2024calibrationcontinuallearningmodels, hwang2025tciltemperaturescalingusing] aims to calibrate CL models for more reliable predictions, whereas our work goes further by proposing a method that enhances calibration and also mitigates forgetting, and improves accuracy. We also introduce new metrics to evaluate calibration and overconfidence in CL, using them to assess the reliability of NC-based methods.
@@
    We provide theoretical analyses to explain why Slerp is more suitable than linear interpolation for generating mixed prototypes (see Sec. 3.3) and to compare different NC-based plasticity losses for handling mixed samples (see Sec. 3.4, Appendix A), with empirical validation of our findings.
@@ 2.1 Continual Learning
CL algorithms can be categorized into two main paradigms: pretrained-based and training-from-scratch. Pretrained methods [wang2022learningpromptcontinuallearning, wang2022dualpromptcomplementarypromptingrehearsalfree, smith2023codapromptcontinualdecomposedattentionbased, Liu_2025_CVPR] leverage large datasets for general representations, while training-from-scratch approaches suit resource-limited settings. These latter algorithms can be grouped into: Architecture-based methods [rusu_2016, yoon_2018, li_learn_tg_2019], which allocate task-specific parameters or expand networks; Regularization-based methods [friedemann_si_2017, sangwon_neurips_2020, fini2022cassle, dongmin, tang_kaizen_wacv_2024, anh_wacv_2025], which constrain weight updates to preserve past knowledge; and Rehearsal-based or replay-based methods [madaan_2022_lump, buzzega2020darkexperiencegeneralcontinual, co2l, wen2024provablecontrastivecontinuallearning, cromomix_mushtaq_eccv_2024], which replay stored data or use generative models.
A primary challenge in CL is the plasticity-stability dilemma - balancing the integration of new knowledge (plasticity) with the retention of past information (stability). Regularization-based methods often address this via knowledge distillation [hinton_kd_2015], which preserves prior knowledge during training. This may involve intermediate features [hou_learning_2019] or final outputs [fini2022cassle, pseudo_negative_icml_2024]. Recent advances extend distillation to relational knowledge, such as instance-wise relation distillation (IRD) in Co2L [co2l] and prototype-instance relation distillation (PRD) in CCLIS [li_cl_important_sampling_2024]. Building on these, [anh_wacv_2025] proposes hardness-softness distillation (HSD), combining IRD and a variant of PRD for fixed prototypes, achieving superior performance. In this work, we adopt HSD to preserve past knowledge through relational consistency with fixed prototypes.
@@ 2.2 Mixup
Mixup [mixup_2018] is a data augmentation strategy based on the principle of vicinal risk minimization [vrm_nips_2000], which interpolates linearly between example pairs and their corresponding labels. The equation of mixup is defined as:
@@
where $(\mathbf{x}_{i},\mathbf{y}_{i})$, $(\mathbf{x}_{j},\mathbf{y}_{j})$ are two random pairs of samples and one-hot labels drawn from the training data, and $\lambda\in[0,1]\sim\text{Beta}(\alpha,\alpha)$, $\alpha\in(0,\infty)$. With simplicity and ease of deployment, mixup has emerged as a well-known method, used widely in both supervised learning [guo2018mixuplocallylinearoutofmanifold, verma2019manifoldmixupbetterrepresentations] and unsupervised learning [kim2020mixcomixupcontrastivelearning, ren2022simpledatamixingprior, shen2022unmixrethinkingimagemixtures] to prevent overfitting and to improve generalization. Moreover, [thulasidasan_on_mixup_training_2020, noh2023rankmixuprankingbasedmixuptraining] show that mixup-trained DNNs are significantly better calibrated, with softmax scores closely reflecting actual correctness likelihood.
Mixup in CL. Mixup has recently been used in CL by combining current task samples with memory instances to promote linearity across tasks. Most studies focus on self-supervised settings [madaan_2022_lump, cromomix_mushtaq_eccv_2024], where mixup serves both as an augmentation technique to increase the number of training samples, especially old ones in memory, and as a regularizer to distill knowledge and reduce forgetting. However, recent work [Maro_as_2021] shows that mixup does not universally guarantee improved calibration and may even degrade it when linear interpolation fails to reflect the underlying data geometry or class proportion. In CL, such conditions naturally arise due to task-wise distribution shifts, imbalance between old and newly introduced classes, and representation drift across sequential updates. Therefore, naively applying linear mixup in CL may worsen confidence-accuracy alignment. These observations motivate a geometry-aware mixing strategy tailored for CL.
Remarks. Like traditional mixup and its CL adaptations, we generates new samples via linear interpolation. However, instead of mixing labels, we interpolate class prototypes on a hypersphere guided by NC geometry. While mixup benefits in CL, its impact on calibration remains underexplored. This work bridges this gap by showing how our method improves calibration, reduces overconfidence, and still boosts accuracy.
Notably, SAMix aims to learn invariant representations across tasks by adopting a two-stage learning framework, where representation learning is decoupled from classifier training. The classifier is used only for evaluation, rather than being jointly optimized with the feature extractor as in one-stage learning. Therefore, existing soft-label mixup methods (e.g., standard mixup [mixup_2018], ReMix [remix_2021]) are not directly applicable, as they mainly affect classifier optimization rather than representation learning. In addition, post-hoc calibration techniques (e.g., label smoothing [szegedy2015rethinkinginceptionarchitecturecomputer] or temperature scaling [guo2017_calibration_nn_2017]) focus only on adjusting prediction confidence and therefore are not compared in this work. Other mix-based augmentations (e.g., CutMix [yun2019cutmixregularizationstrategytrain], Manifold Mixup [verma2019manifoldmixupbetterrepresentations]) operate at the input or feature level and rely on joint classifier optimization. Since SAMix performs geometry-aware mixing at the prototype level to shape representation structure, these methods are not directly comparable to our setting.
@@ 2.3 Neural Collapse
Neural collapse (NC) [papyan_2020] is a phenomenon observed during the final stage of DNN training, where last-layer features and classifier form a structured geometric pattern on a balanced dataset, aligning with a simplex Equiangular Tight Frame (ETF). Building on this foundation, subsequent studies have proven NC is globally optimal in balanced training with cross-entropy [lu2021neuralcollapsecrossentropyloss, zhu_neurips_2021, wenlong_ji_iclr_2022] and mean squared error [poggio_explicit_ra_2020, han_nc_2021, zhou_icml_2022] losses. In addition, [yang_induce_nc_2022] try to induce NC by fixing the classifier under imbalanced training, addressing the issue of misalignment between features and classifier.
The concept of NC is defined as follows.
where $\delta_{i,j}=1$ in case of $i=j$, and 0 otherwise.
Subsequently, the NC phenomenon can be formally characterized by the following four attributes [papyan_2020]:
[…]
These NC attributes provide a conceptual foundation for CL. NC1-NC2 encourage compact and well-separated class representations, which help reduce interference as new classes are introduced. NC3 emphasizes the alignment between feature representations and classifier weights, promoting geometric consistency under sequential learning. As implied by NC4, such alignment induces a favorable decision geometry that simplifies classifier learning and stabilizes adaptation. Together, these properties motivate NC-based approaches for CL, which we build upon in the following section.
NC and Mixup. A recent study [fisher2024pushingboundariesmixupsinfluence] shows that mixup induces a distinct geometric structure. In cross-entropy classifiers with one-hot labels, same-class mixed samples align with a simplex ETF, while cross-class ones form smooth channels along decision boundaries, leading to smoother separation. This effect persists even when classifier weights are fixed as a simplex ETF, enhancing both generalization and calibration.
Inducing NC for CL. Inspired by [galanti_ontr_2022], which showed that NC persists under transfer to new samples or classes, many CL studies [yang2023neuralcollapse, yang2023neuralcollapseterminusunified] leverage NC via DR loss [yang_induce_nc_2022] to reduce forgetting. These methods fix a simplex ETF as classifier prototypes across tasks and align them with sample features. Meanwhile, contrastive learning has gained traction in CL, with remarkable performance from models like Co2L [co2l], CILA [wen2024provablecontrastivecontinuallearning], and CCLIS [li_cl_important_sampling_2024]. [anh_wacv_2025] introduces FNC2, a contrastive loss that explores NC in CL. Here, we employ both FNC2 and DR losses in our analysis.
@@ 3.2 Overall Architecture
Our method is illustrated in Figure 1. We adopt a two-stage learning framework that enhances invariant representations across tasks. As in prior work [co2l, wen2024provablecontrastivecontinuallearning, anh_wacv_2025], we incorporate two main losses: (1) plasticity loss $\mathcal{L}_{plas}$ to learn new knowledge and (2) stability loss $\mathcal{L}_{stab}$ to distill past information. We predefine a set of fixed, non-learnable class prototypes $\mathbf{P}=\{\mathbf{p_{i}}\}_{i=1}^{K}$, structured as a simplex ETF. These prototypes are shared across tasks and are used in both loss terms during training. To encourage smoother and more linear transitions between tasks, we introduce SAMix, an adaptive mixup strategy tailored for NC-based approaches. In replay settings, a memory buffer $\mathcal{M}$ is maintained using reservoir sampling [reservoir_sampling_1985] to store past samples.
The overview of our method is summarized in Algorithm 1. During training, each batch consists of samples from the current dataset $\mathcal{D}_{t}$ and the buffer $\mathcal{M}$. These are augmented and randomly selected and mixed along with their corresponding prototypes (see steps (1)-(2) in Figure 1). All views (original and mixed) are then fed into the current model $f^{t}$ for computing $\mathcal{L}_{plas}$ (steps (3)-(4) in Figure 1), while only original samples are passed through the frozen previous model $f^{t-1}$ to compute the stability loss $\mathcal{L}_{stab}$ (steps (5)–(6) in Figure 1). Mixed prototypes are used solely for $\mathcal{L}_{plas}$ while predefined fixed prototypes are involved in both losses.
Algorithm 1 Sphere-Adaptive Mixup (SAMix) in NC-based Continual Learning. The key SAMix components are the generation of mixed samples and NC-based prototypes via Slerp and their selective use only in the plasticity loss.
3:for $t=1,\cdots,T$ do
4:  for batch $\mathcal{B}_{t}\sim\mathcal{D}_{t}\cup\mathcal{M}$ do
5:   $\lambda\leftarrow\text{Beta}(\alpha,\alpha)$   $\triangleright$ Sample and prototype mixing
7:   $\mathbf{P}_{mix}\leftarrow\text{SAMix}(\mathbf{P},\lambda)$ with Eq. (8)   $\triangleright$ Plasticity loss using both original and mixed samples
8:   if ploss_type is DR then
10:   else
12:   end if  $\triangleright$ Stability loss using only original samples
13:   if t >> 1 then
15:   end if
16:   $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$
17:  end for
19:  $\theta_{prev}\leftarrow\theta$
20:end for